OpenStudy (anonymous):

math help please Solve the following system of equations graphically. y - 4 = 0 2x - y - 2 = 0 What is the solution set?

OpenStudy (anonymous):

a. {(-3, -4)} b. {(-3, 4)} c. {(3, 4)}

OpenStudy (anonymous):

i think its b. but im not sure

OpenStudy (triciaal):

to solve graphically plot the points, draw the lines and find the intersection

OpenStudy (triciaal):

for each line you need 2 points the easiest to use are the value of y when x =0 and the value of x when y = 0

OpenStudy (triciaal):

line 1 y - 4 = 0 so y = 4 line 2 2x - y - 2 = 0 when x = 0 0 - y -2= 0 so y = -2 and when y = 0 2x - 0 -2 =0 so 2x = 2 therefore x = 1 for line 2 the points are (0, -2) and (1, 0 )

OpenStudy (triciaal):

can you sketch the graph and find the intersection?